Sally owns real property for which the annual property taxes are $13,030. She sells the property to Kate on April 2, 2017, for $651,500. Kate pays the real property taxes for the entire year on October 1, 2017. Assume a 365-day year. Round any division to four decimal places. Round your final answers to the nearest dollar.

a. How much of the property taxes can be deducted by Sally and how much by Kate? Sally can deduct $ ------------- and Kate can deduct $ -------------- of the property taxes.

b What effect does the property tax apportionment have on Kate's adjusted basis in the property? Kate's adjusted basis for the property is increased by the $ --------------- she paid that is apportioned to Sally.

c.What effect does the apportionment have on Sally's amount realized from the sale? Sally paid none of the real property taxes and is permitted to deduct the apportioned share of $--------------- . Her amount realized is increased by this amount.

d. How would the answers in parts (b) and (c) differ if the taxes were paid by Sally? If the taxes were paid by Sally, Sally's amount realized would be $---------- . Kate's adjusted basis would be $ .------------
Answer #1

(a) How much of the property taxes can be deducted by Sally and how much by Kate?:-

      Sally = 3/12 * 13030 = $3258 Sally can deduct this amount

      Kate = 9/12 * 13030 = $9772 Kate can deduct this amount

(b) What effect does the property tax apportionment have on Kate's adjusted basis in the property?:-

     Cost 651500 + 3258 Sallys amt she did not pay

      Adjusted basis $654758

(c) What effect does the apportionment have on Sally's amount realized from the sale?

     Selling Price $651500 + 3258 amount realized $654758 because Sally did not pay the property tax & she deduct it.

(d) How would the answers in parts (b) and (c) differ if the taxes were paid by Sally?

Sallys amt realized would be her selling price minus the real property taxes apportioned to Kate ($651500 - $9772 = $641728)

Kate adjusted basis would be her cost minus the real property taxes apportioned to her ($651500 - $9772 = $641728)